Ask your own question, for FREE!
Mathematics 9 Online
OpenStudy (anonymous):

probability Jeanie is a bit forgetful, and if she doesn't make a "to do" list, the probability that she forgets something she is supposed to do is .2. Tomorrow she intends to run three errands, and she fails to write them on her list . What is the probability that Jeanie remembers the first errand but not the second or third?

OpenStudy (anonymous):

4 out of 5 options fulfill this condition for the 1st item of those 4, 1/5 fulfill the second and of the remaining 1/5 fulfill the third. so your answer is found by multiplying (4/5)(1/5)(1/5)

OpenStudy (anonymous):

0.8 * 0.2* 0.2 =4/125

OpenStudy (anonymous):

then viktor.ray isnt right either so dont try that either

OpenStudy (anonymous):

sorry, everything is all right

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

o.O phew

OpenStudy (anonymous):

thank you both)

OpenStudy (anonymous):

it is like if she remember first that means this probability would be 0.8 (1-0.2=0.8), so i need multiply 08*0.2*0.2

OpenStudy (anonymous):

yes exactly

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!